# Opara: Exploiting Operator Parallelism for Expediting DNN Inference on GPUs

### v1

@@ Abstract
GPUs have become the defacto hardware devices to accelerate Deep Neural Network (DNN) inference in deep learning (DL) frameworks. However, the conventional sequential execution mode of DNN operators in mainstream DL frameworks cannot fully utilize GPU resources, due to the increasing complexity of DNN model structures and the progressively smaller computational sizes of DNN operators. Moreover, the inadequate operator launch order in parallelized execution scenarios can lead to GPU resource wastage and unexpected performance interference among operators. To address such performance issues above, we propose Opara, a resource- and interference-aware DNN Operator parallel scheduling framework to accelerate the execution of DNN inference on GPUs. Specifically, Opara first employs CUDA Streams and CUDA Graph to automatically parallelize the execution of multiple DNN operators. It further leverages the resource demands of DNN operators to judiciously adjust the operator launch order on GPUs by overlapping the execution of compute-intensive and memory-intensive operators, so as to expedite DNN inference. We implement and open source a prototype of Opara based on PyTorch in a non-intrusive manner. Extensive prototype experiments with representative DNN and Transformer-based models demonstrate that Opara outperforms the default sequential CUDA Graph in PyTorch and the state-of-the-art DNN operator parallelism systems by up to 1.68×1.68$\times$1.68 × and 1.29×1.29$\times$1.29 ×, respectively, yet with acceptable runtime overhead.
@@ 1 Introduction
Cloud-based GPUs are equipped with an increasing amount of computational power, which typically exceeds the resource demands of individual inference tasks, leading to under-utilization and wastage of hardware resources [4]. To achieve the objective model accuracy with fewer computations, several recent works (e.g., Szegedy et al. [5]) focus on substituting large operators with several smaller and multiple-branch operators in DNN models, which further exacerbates the under-utilization of GPU resources. While batching requests [6] or concurrent processing of multiple model inference tasks [7] can mitigate such GPU under-utilization, it inevitably causes increased latency for model inference due to batching latency and performance interference [8]. As a result, it is compelling to improve the GPU utilization without impacting the DNN inference latency. As DNN models can typically be represented by a Directed Acyclic Graph (DAG) with parallel operators, it provides us an opportunity to exploit operator parallelism for accelerating DNN inference on GPUs while improving the GPU utilization.
Unfortunately, it is nontrivial to efficiently parallelize the execution of DNN operators for a DNN inference task due to the following two facts. First, the model DAG typically exhibits considerable complexity, often incorporating hundreds of operators with complex inter-operator dependencies. For simplicity, existing deep learning (DL) frameworks execute DNN operators one by one in topological sorting order [9], which overlooks the parallelization opportunities among operators. To achieve operator parallelism, a recent work (i.e., Nimble [10]) relies on a reduction transformation of the DNN computation graph, which inevitably brings heavy computation overhead. Second, inadequate operator parallel scheduling can adversely impact the DNN inference performance. As evidenced by motivation experiments in Sec. 2.3, the inadequate operator launch order in mainstream DL frameworks (e.g., PyTorch) can prolong the DNN inference latency by up to $29\%$, due to the GPU blocking caused by the non-preemption feature of CUDA kernels [11] and performance interference among parallelizble operators [8]. In addition, several existing works (e.g., IOS [12]) fail to consider the operator launch and function call overhead due to excessive CPU-GPU interactions when parallelizing DNN operators in the DL framework.
@@ 2.1 DNN Operator Executions on NVIDIA GPUs
To enable parallel executions of operators, we launch operators on multiple CUDA Streams. Each CUDA Stream is actually a task queue that executes tasks sequentially. The execution order of kernels in different CUDA Streams is determined by their arrival order at the stream head. In general, the kernel execution time is considerably short as the batch size is typically small (i.e., ranging from $1$ to $16$) in DNN inference scenarios. Accordingly, the kernel launch overhead constitutes the primary time cost for DNN inference, which offsets the performance gains achieved by the parallel executions of kernels in multiple CUDA Streams. To reduce such overhead, CUDA Graph is a key feature introduced from CUDA 10 that allows scheduling multiple DNN operators on a GPU device at a time.
@@ 2.2 Low GPU Utilization Due to Sequential Execution of DNN Operators
Mainstream DL frameworks generally do not support inter-operator parallel executions for DNN inference, due to the complexity of parallel programming. Instead, they execute DNN operators sequentially in topological sorting order, which cannot fully utilize the GPU resource in general. To illustrate that, we conduct motivation experiments using the stock PyTorch 2.0 with GoogLeNet [14], Inception-v3 [5], and BERT [15] on an NVIDIA A100-PCIE-40GB GPU and an NVIDIA RTX 2080 SUPER GPU. In particular, we adopt the SM efficiency [^1] The SM efficiency denotes the proportion of time during which at least one warp is active and executing instructions on a specific SM in a GPU. measured using NVIDIA Nsight Compute CLI [^2] Nsight Compute CLI: https://docs.nvidia.com/nsight-compute/NsightComputeCli/index.html to evaluate the GPU utilization.
Figure 1: Average SM efficiency of an A100-PCIE-40GB GPU when running GoogLeNet, Inception-v3, and BERT models with different batch sizes.
Figure 1: Average SM efficiency of an A100-PCIE-40GB GPU when running GoogLeNet, Inception-v3, and BERT models with different batch sizes.
Figure 2: Inference latency of GoogLeNet running on an RTX 2080 Super GPU with different operator launch orders and batch sizes.
As shown in Fig. 2, DNN inference on the mainstream DL framework (i.e., the latest PyTorch 2.0) leads to relatively low GPU utilization on A100. Specifically, the SM efficiency of GoogLeNet with the batch size set as $1$ is a mere $2.53\%$, while that of Inception-v3 is $12.04\%$. Even when the batch size increases to $16$, the SM efficiency of GoogLeNet is still $9.12\%$ and that of Inception-v3 reaches $21.48\%$. Similarly, larger models such as BERT cannot fully utilize the GPU resource. By setting the sequence length as $32$, the SM efficiency of BERT with the batch size set as $1$ and $16$ is $18.5\%$ and $54.8\%$, respectively. Meanwhile, we repeat our experiments on a less powerful GPU (i.e., RTX 2080 SUPER), and the SM efficiency of the three workloads ranges from $10.47\%$ to $82.98\%$. Our experiment results indicate that the sequential execution of DNN operators is the root cause of low GPU utilization for running DNN inference, and exploiting the operator parallelism can expedite the DNN inference on GPUs while improving GPU utilization.
@@ 2.3 Performance Impacts of Operator Launch Order
Summary. Low GPU utilization of DNN inference is mainly caused by two factors: First, the sequential execution of DNN operators cannot fully utilize the GPU resources. Second, the default topological sorting order of operator launch is commonly resource- and interference-unaware. Accordingly, judiciously parallelizing the DNN operators with an adequate operator launch order can accelerate DNN inference on GPUs while improving the GPU utilization.
@@ 3.1 Stream Allocator
Problem Formulation and Analysis. As a maximum of $|\mathcal{V}|$ streams can be launched for a DAG, we simply use a matrix $\mathcal{A}$ of size $|\mathcal{V}|\times|\mathcal{V}|$ to represent the stream allocation plan. Each element $a_{ij}\in\mathcal{A}$ is a boolean value, indicating whether the $i$-th operator is executed in the $j$-th stream. Accordingly, a stream allocation plan $\mathcal{A}$ can determine how the DNN operators are parallelized and synchronized. The inference latency $T_{inf}$ of a DNN model is formulated as
@@
where $T_{para}$ denotes the parallelized execution time of a DNN model and $T_{overhead}$ denotes the operator synchronization overhead given a stream allocation plan. In more detail, as the stream allocation plan $\mathcal{A}$ can parallelize the execution of DNN operators, $T_{para}$ can be formulated as
@@
where $T_{seq}$ denotes the inference latency of sequential execution of DNN operators and $h(\mathcal{A})\in(0,1]$ denotes the inference acceleration factor achieved by operator parallelism with the stream allocation plan $\mathcal{A}$. In particular, a certain number of synchronization operators need to be inserted into the model DAG, to ensure the parallelized execution of the model. Such a process introduces the operator synchronization overhead $T_{overhead}$, which is formulated as
@@
0:  DNN computation graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
0:  Set of streams to be launched $\mathcal{S}$ used by the operator set $\mathcal{V}$.
1:  Initialize: $\mathcal{S}\leftarrow\emptyset$;
2:  for each operator $v\in\mathcal{V}$ do
3:     for each predecessor $p\in\mathcal{N}_{pred}$ of $v$ do
4:        if $v$ is the first successor of $p$ then
5:           stream of $v\leftarrow$ stream of $p$;  // put $v$ and $p$ in the same stream
6:           break out of the loop;
7:        end if
8:     end for
9:     if $stream$ of $v$ is $null$ then
10:        $stream$ of $v\leftarrow$ launching a stream;  // put $v$ in a newly launched stream
11:        $\mathcal{S}\leftarrow\mathcal{S}\cup\left\{stream\text{ of}\>v\right\}$;
12:     end if
13:  end for
14:  return  $\mathcal{S}$.
@@
Stream Allocation Algorithm. To solve such a complex stream allocation problem in polynomial time, we design a heuristic algorithm in Alg. 1. The key idea of our algorithm is to allocate parallelizable operators to multiple CUDA Streams as much as possible (i.e., minimizing the value of $h(\mathcal{A})$). To avoid excessive synchronization operations (i.e., minimizing the value of $g(\mathcal{A})$), we aim to greedily put non-root nodes (i.e., operators) in the same CUDA stream as one of their predecessor operators. Specifically, given a computation graph $\mathcal{G}$, Opara first initializes a set of streams to be launched $\mathcal{S}$ and then enumerates operators in $\mathcal{V}$ in topological sorting order (lines $1$-$2$). For each operator ${v}\in\mathcal{V}$, it iterates over all of its predecessors $p\in\mathcal{N}_{pred}$ (line $3$). If $v$ is the first successor of the current predecessor $p$, it allocates $v$ to the same stream of $p$; otherwise, it moves on to the next predecessor (lines $4$-$7$). If $v$ does not find a predecessor that satisfies such a condition above, we allocate the operator $v$ to a newly launched stream (lines $9$-$11$). In particular, the streams consume GPU resources only when operators are running, which indicates that performance interference among streams does not occur when operators are not executed on GPUs.
@@ 3.3 Operator Launcher
Problem Analysis. As illustrated in Sec. 2.3, inadequate operator launch orders can significantly affect the DNN inference latency. To identify an optimal launch order, a naive solution is iterating through all possible topological sorting orders of a model DAG and choose the order with the lowest inference latency. However, such a method involves selecting nodes with zero indegree and deleting the corresponding vertices and their connected edges. By assuming that $n$ operators exist in a model DAG, the time complexity of traversing all topological sorting orders is $\mathcal{O}(n!)$, which is also an NP-hard problem. As a result, we turn to designing a heuristic operator launch algorithm to solve such a complex problem.
0:  DNN computation graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
0:  Operator queue $\mathcal{Q}$ in resource- and interference-aware operator launch order.
1:  Initialize: List of operators to be launched $\mathcal{L}\leftarrow\emptyset$, $\mathcal{L}_{mem}\leftarrow\emptyset$, $\mathcal{L}_{comp}\leftarrow\emptyset$, and $\mathcal{Q}\leftarrow\emptyset$;
2:  Add the operators $v\in\mathcal{V}$ with an indegree of 00 that is memory-intensive and compute-intensive to $\mathcal{L}_{mem}$ and $\mathcal{L}_{comp}$, respectively;
3:  while $\mathcal{L}_{mem}$ or $\mathcal{L}_{comp}$ is not empty do
4:     $\mathcal{L}$ = alternately choose a non-empty list from ${\{\mathcal{L}_{mem},\mathcal{L}_{comp}\}}$;
5:     $v_{min}\leftarrow$ the operator that requires the least amount of GPU resources in $\mathcal{L}$;
6:     .$\mathcal{L}.{remove}({v_{min}})$, and $\mathcal{Q}.{append}({v_{min}})$;   // launch the operator $v_{min}$
7:     for each successor $s\in\mathcal{N}_{succ}$ of $v_{min}$ do
8:        $indegree$ of ${s}\leftarrow$ $indegree$ of ${s}-1$;   // update the $indegree$ of ${s}$
9:        if $indegree$ of ${s}==0$ then
10:           if operator ${s}$ is memory-intensive then
11:              $\mathcal{L}_{mem}.{append}({s})$;  // add $s$ to $\mathcal{L}_{mem}$
12:           else
13:              $\mathcal{L}_{comp}.{append}({s})$;  // add $s$ to $\mathcal{L}_{comp}$
14:           end if
15:        end if
16:     end for
17:  end while
18:  return  $\mathcal{Q}$.

### v2

@@ Abstract
GPUs have become the defacto hardware devices for accelerating Deep Neural Network (DNN) inference workloads. However, the conventional sequential execution mode of DNN operators in mainstream deep learning frameworks cannot fully utilize GPU resources, even with the operator fusion enabled, due to the increasing complexity of model structures and a greater diversity of operators. Moreover, the inadequate operator launch order in parallelized execution scenarios can lead to GPU resource wastage and unexpected performance interference among operators. In this paper, we propose Opara, a resource- and interference-aware DNN Operator parallel scheduling framework to accelerate DNN inference on GPUs. Specifically, Opara first employs CUDA Streams and CUDA Graph to parallelize the execution of multiple operators automatically. To further expedite DNN inference, Opara leverages the resource demands of operators to judiciously adjust the operator launch order on GPUs, overlapping the execution of compute-intensive and memory-intensive operators. We implement and open source a prototype of Opara based on PyTorch in a non-intrusive manner. Extensive prototype experiments with representative DNN and Transformer-based models demonstrate that Opara outperforms the default sequential CUDA Graph in PyTorch and the state-of-the-art operator parallelism systems by up to 1.68×1.68$\times$1.68 × and 1.29×1.29$\times$1.29 ×, respectively, yet with acceptable runtime overhead.
@@ 1 Introduction
Cloud-based GPUs are equipped with an increasing amount of computational power, which typically exceeds the resource demands of individual inference tasks, leading to under-utilization and wastage of hardware resources [4]. To achieve the objective model accuracy with fewer computations, several recent works (e.g., Szegedy et al. [5]) focus on substituting large operators with several smaller and multiple-branch operators in DNN models, which further exacerbates the issue of GPU under-utilization. While batching requests [6] or co-locating model inference tasks [7] on a GPU can mitigate such GPU under-utilization, it inevitably prolongs the model inference due to batching latency and performance interference [8]. Moreover, the operator fusion [9] cannot fully utilize the GPU resources (as discussed in Sec. 2.2) due to the limited scope of pre-defined fusion rules. Fortunately, as DNN models can typically be represented by a Directed Acyclic Graph (DAG) with parallel operators, it provides us an opportunity to exploit inter-operator parallelism for accelerating DNN inference on GPUs while improving the GPU utilization.
However, it is nontrivial to efficiently parallelize the execution of DNN operators for a DNN inference task due to the following two facts. First, the model DAG typically exhibits considerable complexity, often incorporating hundreds of operators with complex inter-operator dependencies. For simplicity, existing deep learning (DL) frameworks execute DNN operators one by one in topological sorting order [10]. To achieve operator parallelism, a recent work (i.e., Nimble [11]) relies on a reduction transformation of the DNN computation graph, which inevitably brings heavy computation overhead. Second, inadequate operator parallel scheduling can adversely impact the DNN inference performance. As evidenced by motivation experiments in Sec. 2.3, the inadequate operator launch order in mainstream DL frameworks (e.g., PyTorch) can prolong the DNN inference latency by up to $29\%$, due to the GPU blocking caused by the non-preemption feature of CUDA kernels [12] and performance interference among parallelizable operators [8]. In addition, several existing works (e.g., IOS [13]) fail to consider the operator launch and function call overhead due to excessive CPU-GPU interactions when parallelizing DNN operators in the DL framework.
@@ 2.1 DNN Operator Executions on NVIDIA GPUs
To enable parallel executions of operators, we launch operators on multiple CUDA Streams. Each stream is actually a task queue that executes tasks sequentially. The execution order of kernels in different CUDA Streams is determined by their arrival order at the stream head. In general, the kernel execution time is considerably short as the batch size is typically small (i.e., ranging from $1$ to $16$) in latency-critical inference scenarios [15, 9]. Accordingly, the kernel launch overhead constitutes the primary time cost for DNN inference, which negatively impacts the performance gains achieved by the parallel executions of kernels in multiple CUDA Streams. To reduce such overhead, CUDA Graph is a key feature introduced from CUDA 10 that allows scheduling multiple DNN operators on a GPU device at a time.
@@ 2.2 Low GPU Utilization Due to Sequential Execution of DNN Operators
Mainstream DL frameworks execute DNN operators sequentially in topological sorting order, which cannot fully utilize GPU resources. To illustrate that, we conduct motivation experiments using the stock PyTorch 2.0 and ONNX Runtime 1.12111https://onnxruntime.ai/ with the operator fusion [9] enabled. We serve three typical DNN inference models including GoogLeNet [^2] https://pytorch.org/hub/pytorch_vision_googlenet/, Inception-v3 [5], and BERT [^3] https://huggingface.co/google-bert on both an NVIDIA A100-PCIE-40GB GPU and an NVIDIA RTX 2080 SUPER GPU. In particular, we adopt the SM efficiency measured using NVIDIA Nsight Compute CLI [^4] https://docs.nvidia.com/nsight-compute/NsightComputeCli/index.html to evaluate the GPU utilization.
Figure 1: Average SM efficiency of an NVIDIA A100-PCIE-40GB GPU when running GoogLeNet, Inception-v3, and BERT models in PyTorch and ONNX Runtime.
Figure 2: Inference latency of GoogLeNet running on an NVIDIA RTX 2080 Super GPU with different operator launch orders and batch sizes.
As shown in Fig. 2, DNN inference on the mainstream DL frameworks achieves relatively low to medium GPU utilization even with the operator fusion technique enabled. Specifically, the SM efficiency of GoogLeNet, Inception-v3, and BERT with the batch size as $1$ is merely $2.53\%$, $12.04\%$, and $18.5\%$, respectively, achieved in the stock PyTorch on an A100 GPU. Even when the batch size is increased to $16$ and serving DNN inference in the ONNX Runtime, the SM efficiency of the three workloads is moderately increased to $57.21\%$, $39.9\%$, and $76.37\%$, respectively. Moreover, we repeat our experiments on a less powerful GPU (i.e., RTX 2080 SUPER), and the SM efficiency of the three workloads ranges from $10.47\%$ to $82.98\%$. Such experiment results above indicate that (1) the sequential execution of DNN operators is the root cause of low GPU utilization for serving DNN inference; (2) The operator fusion technique can only combine a certain number of parallelizable operators based on the pre-defined fusion rules [9], resulting in moderate GPU utilization. Accordingly, there still exists enough room to exploit the operator parallelism for improving the GPU utilization of DNN inference on GPUs.
@@ 2.3 Performance Impacts of Operator Launch Order
Summary. Low GPU utilization of DNN inference is mainly caused by two factors: First, the sequential execution of DNN operators cannot fully utilize the GPU resources, even with the operator fusion enabled. Second, the default topological sorting order of operator launch is commonly resource- and interference-unaware. Accordingly, judiciously parallelizing DNN operators with an adequate operator launch order is compelling for accelerating DNN inference on GPUs while improving the GPU utilization.
@@ 3.1 Stream Allocator
Problem formulation and analysis. As a maximum of $|\mathcal{V}|$ streams can be launched for a DAG, we simply use a matrix $\mathcal{A}$ of size $|\mathcal{V}|\times|\mathcal{V}|$ to represent the stream allocation plan, which determines how the DNN operators are parallelized and synchronized. Each element $a_{ij}\in\mathcal{A}$ is a boolean value, indicating whether the $i$-th operator is executed in the $j$-th stream. We formulate the inference latency $T_{inf}$ as
@@
where $T_{para}$ denotes the parallelized execution time of a DNN model and $T_{overhead}$ denotes the operator synchronization overhead given a stream allocation plan $\mathcal{A}$. In more detail, as the plan $\mathcal{A}$ can parallelize the execution of DNN operators, $T_{para}$ can further be formulated as
@@
where $T_{seq}$ denotes the inference latency of sequential execution of DNN operators and $h(\mathcal{A})\in(0,1]$ denotes the inference acceleration factor achieved by operator parallelism with the plan $\mathcal{A}$. To ensure the parallelized execution of the model, a certain number of synchronization operators [16] require to be inserted into the model DAG. Such a process introduces significant operator synchronization overhead $T_{overhead}$, which can be formulated as
@@
Input : DNN computation graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$
Output : Set of streams to be launched $\mathcal{S}$
1 Initialize: $\mathcal{S}\leftarrow\emptyset$, $\mathtt{SYNC}$ $flag$ $\leftarrow\mathtt{False}$ for each operator $v\in\mathcal{V}$, and sort $\mathcal{V}$ in topological sorting order;
2 for each operator $v\in\mathcal{V}$ do
3       for each predecessor $p\in\mathcal{N}_{pred}$ of $v$ do
4             if $flag$ of $p$ is $\mathtt{False}$ then
                   $stream$ of $v\leftarrow$ $stream$ of $p$; // put $v$ and $p$ in the same stream
5                   $flag$ of $p\leftarrow\mathtt{True}$;
6                   break out of the loop;
7
8
9      if $stream$ of $v$ is $null$ then
             $stream$ of $v\leftarrow$ launching a stream; // put $v$ in a newly launched stream
10             $\mathcal{S}\leftarrow\mathcal{S}\cup\left\{stream\text{ of }v\right\}$;
11
12
13return $\mathcal{S}$;
@@ 10             $\mathcal{S}\leftarrow\mathcal{S}\cup\left\{stream\text{ of }v\right\}$;
Stream allocation algorithm. The key idea of Alg. 1 is to allocate parallelizable operators to multiple CUDA Streams as much as possible (i.e., minimizing the value of $h(\mathcal{A})$). Moreover, we greedily put non-root nodes (i.e., operators) in the same CUDA stream as one of their predecessor operators, so as to avoid excessive synchronization operators (i.e., minimizing the value of $g(\mathcal{A})$). Specifically, given a computation graph $\mathcal{G}$, Opara first initializes a set of streams to be launched $\mathcal{S}$ and the $\mathtt{SYNC}$ $flag$ of each $v\in\mathcal{V}$. It then enumerates operators in $\mathcal{V}$ in topological sorting order (lines $1$-$2$). For each operator ${v}\in\mathcal{V}$, it iterates over all of its predecessors $p\in\mathcal{N}_{pred}$ (line $3$). If the current predecessor $p$ has not yet contributed to reducing the synchronization overhead (i.e., $flag$ is $\mathtt{False}$), it allocates $v$ to the same stream of $p$, and set the $flag$ of $p$ as $\mathtt{True}$ (lines $4$-$7$). If $v$ does not find a predecessor that satisfies such a condition above, we allocate the operator $v$ to a newly launched stream (lines $8$-$10$). In particular, the parallelized execution of streams does not impact each other as long as operators are not executed on GPUs. To ease the understanding of Alg. 1, we present an illustrative example in Appendix .1.
@@ 3.3 Operator Launcher
Problem analysis. As illustrated in Sec. 2.3, inadequate operator launch orders can significantly affect the DNN inference latency. To identify an optimal launch order, a naive solution is iterating through all possible topological sorting orders of a model DAG and choosing the order with the lowest inference latency. However, such a method involves selecting nodes with zero indegree and deleting the corresponding vertices and their connected edges. By assuming $n$ operators exist in a model DAG, the time complexity of traversing all topological sorting orders is $\mathcal{O}(n!)$, which is also an NP-hard problem [17]. As a result, we turn to designing a heuristic operator launch algorithm to solve such a complex problem.
Input : DNN computation graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$
Output : Operator queue $\mathcal{Q}$ in resource- and interference-aware operator launch order
1 Initialize: List of operators to be launched $\mathcal{L}\leftarrow\emptyset$, $\mathcal{L}_{mem}\leftarrow\emptyset$, $\mathcal{L}_{comp}\leftarrow\emptyset$, and $\mathcal{Q}\leftarrow\emptyset$;
2 Add the operators $v\in\mathcal{V}$ with an $indegree$ of 00 that are memory-intensive and compute-intensive to $\mathcal{L}_{mem}$ and $\mathcal{L}_{comp}$, respectively;
3 while $\mathcal{L}_{mem}$ or $\mathcal{L}_{comp}$ is not empty do
4       $\mathcal{L}\leftarrow$ alternately choose a non-empty list from ${\{\mathcal{L}_{mem},\mathcal{L}_{comp}\}}$;
5       $v_{min}\leftarrow$ the operator that requires the least amount of GPU resources in $\mathcal{L}$;
       .$\mathcal{L}.{remove}({v_{min}})$, and $\mathcal{Q}.{append}({v_{min}})$; // launch the operator $v_{min}$
6       for each successor $s\in\mathcal{N}_{succ}$ of $v_{min}$ do
7             Update $indegree$ of ${s}\leftarrow$ $indegree$ of ${s}-1$;
8             if $indegree$ of ${s}==0$ then
9                   if operator ${s}$ is memory-intensive then
                         $\mathcal{L}_{mem}.\text{append}(s)$; // add $s$ to $\mathcal{L}_{mem}$
10
11                  else
                         $\mathcal{L}_{comp}.\text{append}(s)$; // add $s$ to $\mathcal{L}_{comp}$
12
13
14
15
16return $\mathcal{Q}$;